### v1

@@ 2.1 Muonium HFS
Currently the muonium hyperfine splitting = 4 463 302 765(53) Hz is determined with an experimental precision of 12 ppb, and the ratio of the muon and proton magnetic moments at 120 ppb from the previous high-field experiment at LAMPF [14] mostly dominated by statistical errors. By combining the world’s most intense pulsed muon beam at the MUSE H-line [15, 16] with a new analysis method “Rabi oscillation spectroscopy” [17] (see below), we aim at measuring and with an order of magnitude higher precision.
@@ 3.2 Measurement strategy
The muonium resonance frequencies and will first be measured in a magnetic field of 1.7 T following the approach of the last experiment [14]. Then, the magnetic field uncertainty can be reduced by measuring at the maximum field strength produced by our MRI magnet (2.9 T) since the magnetic probe accuracy is independent of the applied field. Finally, we plan to measure at a field of 1.15 T, where the effect of the fluctuating magnetic field is minimal because the derivative of and becomes zero. Combining these three measurements, which will require different microwave cavities (see below), will put a stringent test on the internal consistency of the experiment’s systematic uncertainty.
@@ 4.1 Uniform magnetic field and magnetic probes
A continuous wave nuclear magnetic resonance (CW-NMR) system with an RF pickup coil was developed to precisely measure the magnetic field for both /EDM and MuSEUM experiments at J-PARC [43]. The development of the CW-NMR probe is reported in detail in [44]. Figure 5(a) shows the developed standard probe, which recently achieved a magnetic field measurement accuracy of 15 ppb.
@@ 4.2 Microwave system
Muonium transition frequencies in a 1.7 T magnetic field are approximately 1.906 GHz and 2.556 GHz. A cylindrical microwave cavity was already designed and constructed for the high-field measurement to have two frequency modes, TM110 and TM210, to match the resonance frequencies and [46, 47]. The frequency ratio of TM110 and TM210 modes yields a constant value for a cylindrical cavity, independently of the cavity radius. Thus, a cylindrical cavity can only be used at 1.7 T (magic field), where the resonance frequency relation holds. As shown in Fig. 6(a), there are two input ports, one for each frequency, each equipped with a piezo actuator directly attached to the gas chamber for tuning by displacing a conductive tuning bar, covering the full sweep range of the resonance line shape. The microwave cavity shown in Fig. 6(b) is made of oxygen-free copper with windows 25-m thick to let muons in and positrons out. The cavity with an 187 mm inner diameter is 304 mm long as compared to 160 mm for the LAMPF experiment to allow measurement at lower Kr gas density, thus reducing the systematic uncertainty on the gas density extrapolation to zero pressure. The stored microwave power can be monitored through a pickup antenna with a power meter. Since resonance lines are distorted by fluctuating microwave power, the signal from the pickup antenna in the cavity will be used to monitor and regulate the power to achieve a stability of 0.01% [47].
@@ 4.4.4 Silicon strip detector
A new type of positron (electron) detector made of a highly-segmented silicon strip sensor with high-rate capability, which was originally developed for the J-PARC /EDM experiment, was used and tested (Fig. 7(b)) [57]. The silicon strip sensor has an active area of 97.28 mm square divided into two blocks with a thickness of 0.32 mm. The strip pitch and length are 0.19 mm and 48.575 mm, respectively. The number of strips is 512 2 blocks. The silicon strip sensor is connected on both sides through an adapter to two multi-Slit128A boards, where an analog/digital combined type integrated circuit directly processes signals. Each board contains four readout SliT128A chips, ASIC-based ASD, that are controlled by an FPGA. The details and performance of this silicon strip detector are reported in [58].
@@ 4.6 Analysis method
In standard spectroscopy, the resonance curve is drawn by sweeping the microwave frequency and measuring the spin-flip signal to determine the resonance frequency from the center of the resonance curve. In the processor experiment, a resonance line narrowing technique was adopted to improve the accuracy that only selects long-lived muonium atoms (“old muonium” method) [14]. However, since the muon beam provided at LAMPF was quasi-continuous, 70% of the muon flux was discarded to produce a pulsed time structure. Also, the “old muonium” technique was not applied in the analysis but rather implemented in the hardware.
@@ Acknowledgements
The muon experiment at the Materials and Life Science Experimental Facility of J-PARC was performed under a user program (Proposals No. 2020B0333, 2021B0169, 2022A0159, 2022B0314, 2023A0317, 2024A0075).
@@
Funding Information This work was supported by the JSPS KAKENHI Grant No. 26247046, 17H01133, 21H04463, and 21H04481.
Competing Interests (not applicable)
Ethical Standards (not applicable)
Consent for Publication (not applicable)

### v2

@@ 2.1 Muonium HFS
Currently the muonium hyperfine splitting = 4 463 302 765(53) Hz is determined with an experimental precision of 12 ppb, and the ratio of the muon and proton magnetic moments at 120 ppb from the previous high-field experiment at LAMPF [14] mostly dominated by statistical errors. By combining the world’s most intense pulsed muon beam at the MUSE H-line [15, 16] with a new analysis method “Rabi oscillation spectroscopy” [17] (see Sect. 4.6), we aim at measuring and with an order of magnitude higher precision.
@@ 3.2 Measurement strategy
The muonium resonance frequencies and will first be measured in a magnetic field of 1.7 T following the approach of the last experiment [14]. Then, the magnetic field uncertainty can be reduced by measuring at the maximum field strength produced by our MRI magnet (2.9 T) since the magnetic probe accuracy is independent of the applied field. Finally, we plan to measure at a field of 1.15 T, where the effect of the fluctuating magnetic field is minimal because the derivative of and becomes zero. Combining these three measurements, which will require different microwave cavities (see Sect. 4.2), will put a stringent test on the internal consistency of the experiment’s systematic uncertainty.
@@ 4.1 Uniform magnetic field and magnetic probes
A continuous wave nuclear magnetic resonance (CW-NMR) system with an RF pickup coil was developed to precisely measure the magnetic field for both muon /EDM and MuSEUM experiments at J-PARC [43]. The development of the CW-NMR probe is reported in detail in [44]. Figure 5(a) shows the developed standard probe, which recently achieved a magnetic field measurement accuracy of 15 ppb.
@@ 4.2 Microwave system
Muonium transition frequencies in a 1.7 T magnetic field are approximately 1.897 GHz and 2.566 GHz. A cylindrical microwave cavity was already designed and constructed for the high-field measurement to have two frequency modes, TM110 and TM210, to match the resonance frequencies and [46, 47]. The frequency ratio of TM110 and TM210 modes yields a constant value for a cylindrical cavity, independently of the cavity radius. Thus, a cylindrical cavity can only be used at 1.7 T (magic field), where the resonance frequency relation holds. As shown in Fig. 6(a), there are two input ports, one for each frequency, each equipped with a piezo actuator directly attached to the gas chamber for tuning by displacing a conductive tuning bar, covering the full sweep range of the resonance line shape. The microwave cavity shown in Fig. 6(b) is made of oxygen-free copper with windows 25-m thick to let muons in and positrons out. The cavity with an 187 mm inner diameter is 304 mm long as compared to 160 mm for the LAMPF experiment to allow measurement at lower Kr gas density, thus reducing the systematic uncertainty on the gas density extrapolation to zero pressure. The stored microwave power can be monitored through a pickup antenna with a power meter. Since resonance lines are distorted by fluctuating microwave power, the signal from the pickup antenna in the cavity will be used to monitor and regulate the power to achieve a stability of 0.01% [47].
@@ 4.4.4 Silicon strip detector
A new type of positron (electron) detector made of a highly-segmented silicon strip sensor with high-rate capability, which was originally developed for the J-PARC muon /EDM experiment, was used and tested (Fig. 7(b)) [57]. The silicon strip sensor has an active area of 97.28 mm square divided into two blocks with a thickness of 0.32 mm. The strip pitch and length are 0.19 mm and 48.575 mm, respectively. The number of strips is 512 2 blocks. The silicon strip sensor is connected on both sides through an adapter to two multi-Slit128A boards, where an analog/digital combined type integrated circuit directly processes signals. Each board contains four readout SliT128A chips, ASIC-based ASD, that are controlled by an FPGA. The details and performance of this silicon strip detector are reported in [58]. This detector was used in an experiment at zero field in June 2018 to apply Rabi-oscillation spectroscopy to muonium HFS measurements (see Sects. 4.6 and 5, while for the high-field experiment, an improved version with four sensors mounted on one board and new readout ASIC chips SliT128D for better performance will be used.
@@ 4.6 Analysis method
In standard spectroscopy, the resonance curve is drawn by sweeping the microwave frequency and measuring the spin-flip signal to determine the resonance frequency from the center of the resonance curve. In the predecessor experiment, a resonance line narrowing technique was adopted to improve the accuracy that only selects long-lived muonium atoms (“old muonium” method) [14]. However, since the muon beam provided at LAMPF was quasi-continuous, 70% of the muon flux was discarded to produce a pulsed time structure. Also, the “old muonium” technique was not applied in the analysis but rather implemented in the hardware.
@@ Acknowledgements
The muon experiment at the Materials and Life Science Experimental Facility of J-PARC was performed under a user program (Proposals No. 2017A0134, 2017B0224, 2018A0071, 2018B0314, 2019A0144, 2020B004, 2020B0333, 2021B0169, 2022A0159).
@@
Funding Information This work was supported by the JSPS KAKENHI Grant No. 26247046, 17H01133, 19K14746, 21H04463, and 21H04481.